A system absorbs 11.5 KJ of heat while it is compressed from 5.5 L to 1.0 P by a pressure of 7 atm. Find delta U in KJ for the system
Answer #1

The work done on the system (w) = -P.\DeltaV = -P(V2 - V1) = -7 atm * (1 - 5.5) L = 31.5 L.atm = 3189.9 J or 3.19 kJ

According to the definition of enthalpy change, \DeltaH = \DeltaU + (-w), i.e. 11.5 kJ = \DeltaU + (-3.19), i.e. \DeltaU = 14.7 kJ